# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v2

@@ IV.3 Influence of Encoding Circuits
Figure 8: Insight into the impact of data encoding circuits on test performance scores. Results are aggregated across all datasets within the respective family and $n_{\mathrm{layers}}\in[1,8]$. The upper panel exemplifies this for the Friedman regression problem. In (a) we show results corresponding to the dataset with $d=15$ features. Here, test performance scores, as measured by MSE, are aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel (FQK/PQK). The findings in (b) represent test performance scores aggregated over all datasets in the Friedman family. The lower panel details the two curves diff dataset. In (c), we display test performances, as measured by ROC-AUC score, aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with degree $D=13$. Aggregated results across all datasets are depicted in (d).
@@
To investigate the impact of data encoding circuits on the performance of the resulting models, we aggregate results across all datasets within each family and choose $n_{\mathrm{layers}}$ varying from 1–8. We then show the results for each encoding circuit. This is exemplified in Fig. 8 for the Friedman (regression) and the two curves diff (classification) datasets. The results corresponding to the other dataset families of this study are shown in Fig. A12 of the Appendix D.3. When examining the performance scores for a single dataset within a dataset family, significant performance differences among the data encoding circuits emerge. This can be seen in Figs. 8 (a) and (c). Here, the “HighDimEncodingCircuit” and the “ZZFeatureMap” perform particularly poor in regression tasks, while the “ChebyshevPQC encoding circuit” underperforms in classification tasks, both for FQK and PQK approaches (note that trends for MSE and ROC-AUC performance scores work in opposite directions). However, when considering performance scores aggregated across all datasets within a dataset family, we observe that, except for the “ChebyshevPQC” in the QSVC-FQK case and with some trend for the “ZZFeatureMap” in regression tasks, all encoding circuits perform comparably in terms of overall performance and variance (cf. Figs. 8 (b) and (d)). This supports arguments for problem-specific data encoding architecture search [43, 24]. Furthermore, this aggregated analysis does not reveal any significant performance differences between FQK and PQK approaches.
@@
In earlier studies [34], it has been found that data encoding circuits without entanglement can perform surprisingly well. In our case, these are the “SeparableRx” and “ZFeatureMap” encoding circuits. As demonstrated in Ref. [27] for FQKs, the “SeparableRx” circuit is analytically tractable, and this can be straightforwardly extended to PQKs. Similarly, the “ZFeatureMap”, which incorporates additional Hadamard gates, also demonstrates comparable performance. With the model performances obtained in Fig. 8, this raises the possibility that driving forces other than “quantumness” might be fostering the performance of QKMs. This observation underscores the need for further comprehensive and systematic investigations into the underlying factors contributing to QKM performance.
IV.4 Influence of Data Encoding Strategies
As already implicitly indicated above, it is known that the number of qubits and layers of data encoding circuits determine the Fourier frequency spectrum of the resulting quantum model and that encoding features redundantly can help in getting more accurate results [11, 14]. However, from a practical point of view it is not clear how to best distribute the features of a data point among the available qubits in a particular encoding circuit. To shed some light on this aspect, we compare QKMs built from selected encoding circuits following the two encoding schemes illustrated in Fig. 9. In both embedding schemes, features are encoded in the qubits from top to bottom. Differences in the two encoding strategies occur when the number of qubits is larger than the number of features. In “option 1”, after all features have been encoded once in a layer, the encoding is repeated and then cut off once the final qubit is reached. In “option 2”, the enumeration of features is not reset at the next layer. As a consequence, each feature in “option 1” is always assigned to the same qubit, whereas in “option 2” the features can be shuffled across qubits.
Figure 9: Schematic illustration of two different data encoding options considered in the investigations of Figs. 10 and  11.
Using the QFMNIST dataset with $d=5$ as an example, Fig. (10) shows the findings for this regression tasks. The results for each quantum kernel are aggregated across the regression methods and the respective encoding circuits. Then, the median of the resulting MSE test score is calculated. For FQK, both embedding options show that largest scores are achieved for ten qubits, i.e., for integer multiples of the number of features. It can be seen that higher scores occur more frequently with embedding “option 2”. For PQK, it appears that no feature redundancies are necessary in either case. Here, however, embedding “option 1” often delivers larger scores. Overall, the MSE differences are small.
Figure 10: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the QFMNIST regression dataset with $d=5$ principal components. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
Figure 11 displays the results for the two curves diff classification dataset for $D=13$ (the number of features is still fixed to $d=4$). Again, results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is evaluated. It can be seen that here in contrast to the previous example, both embedding options deliver the best classification scores for non-integer multiples of the number of features (for PQK and FQK). Here, “option 1” performs better in each case. In addition, larger values of $n_{\text{layers}}$ and $n_{\text{qubits}}$ lead to increasing classification scores. The difference to the previous example (Fig. 10) might be due to significantly larger dataset complexity. This also explains the overall larger difference in performance scores for the two curves diff .
Figure 11: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK- and PQK-QSVC approaches, respectively for the two curves diff classification dataset with $D=13$. Note, the feature dimension is kept constant at $d=4$. The results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
The trends discussed here also apply to the other datasets in this study. These are summarized in detail in Appendix D.4.
IV.5 Analysis of PQK Design Options
Projected quantum kernels as defined in Eq. (8) have additional degrees of freedom which can strongly influence the performance of the resulting model. Specifically, any proper outer kernel function $\kappa$ can be applied and the measurement operator $O^{k}$ can be chosen freely. To the best of our knowledge, a thorough investigation of the influence of both of these design choices on the model is still lacking. [^4] Reference [37] touches upon similar aspects but has a much narrower focus on the impact of the RDM size on cross-validation accuracy and generalization ability.
[…]
We aim to quantify the influence of the respective choices on the model performance, while taking into account the corresponding impact of all other hyperparameters. For this purpose, we aggregate all simulation results per dataset and conduct (semi-) partial correlation analyses (cf. Appendix. A for details). The results are summarized in Tab. 1. Note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, the opposite is the case for correlations with ROC-AUC scores. An overview of the model performances used for calculating the values in the table is provided in Appendix D.5 in Figs. A16 and A17.
@@ IV.5 Analysis of PQK Design Options
Table 1: Overview of semi-partial correlation analysis results for exploring the effect of the choice of the outer kernel function $\kappa$ (cf. (8)) and the measurement operator $O^{k}$ (cf. Eqs. (14), (16), (18)) in the definition of a PQK on the resulting test performance scores, while controlling for the effect of all other hyperparameters on the test performance. We investigate PQK-QSVC and PQK-QSVR models for classification and regression tasks, respectively. We compute ROC-AUC scores for evaluating classification accuracy and MSEs for assessing regression results. We employ Spearman $\rho$ rank-order correlation analysis as implemented in the pingouin [59] package. Bold values refer to statistically significant correlations, with $p$-values0.05absent0.05$\leq$ 0.05≤ 0.05, while others refer to not statistically significant correlation, which are given for the sake of completeness. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
@@
From the table we directly realize that both $\kappa$ and the choice of $O^{k}$ for defining the respective projected quantum circuits exhibit statistically significant correlations. For the outer kernel, no universal statement can be made a priori regarding which one is generally suitable for effective QKM design—at least not within the scope of our study. The only exception is that for the classification tasks considered here, the RationalQuadratic kernel is not a good choice. Furthermore, it is worth noting that the commonly used vanilla PQK in the literature, i.e., with a Gaussian kernel, is not automatically the best choice.
@@
Regarding a proper choice of the measurement operator $O^{k}$, Tab. 1 suggests $X_{k=1}$ to be insufficient for all regression and classification tasks considered here. In this case, we frequently observed training kernel Gram matrices entirely filled with ones, thus rendering them impractical for learning. In contrast, utilizing the $P^{1+2}$ measurement operator, cf. Eq. (18), points towards an overall improvement in model performance (within the scope of this study).
Apart from that, the results indicate that the definition of the projected quantum circuits via $O^{k}$, as well as the choice of $\kappa$ should be considered as additional hyperparameters for PQK model design. A systematic investigation of this is however still lacking.
@@
In Ref. [34] the authors have raised the question whether the projected quantum circuit is not so much responsible for learning, but rather the subsequent outer (Gaussian) kernel applied to the features computed by Eq. (II). Referring to Tab. 1 per dataset, it becomes evident that both the choice of $O^{k}$ (and thus the projected quantum circuit itself), and the outer kernel $\kappa$ are crucial for achieving sufficiently good model performance. Therefore, the question of which factor is more important cannot be definitively answered based on this analysis alone. Consequently, we further list all (statistically significant) correlations of the various hyperparameters of the PQK QKMs for a given outer kernel, both among themselves and with model performance in Appendix D.5 in Tab. 2. Thereof, it again becomes clear that depending on the dataset and outer kernel function, both hyperparameters that contribute to the definition of the projected quantum circuits (i.e., $n_{\mathrm{layers}}$ and $w_{\mathrm{e}}$) and those solely associated with the outer kernel exhibit statistically significant correlations.
@@
where $|G|$ refers to the number of entries in $G$. For Gaussian outer kernels it directly follows from Eq. (9) that the projected quantum circuit contribution $F_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ becomes
$F_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=-\frac{\log[k_{\mathbf{\theta}}^{\text{PQK}}(\mathbf{x},\mathbf{x}^{\prime})]}{\gamma}\,.$ (20)
Thus, calculating $d(G|G^{\prime})$ also defines a difference measure between projected quantum circuit contributions $F$ and $F^{\prime}$, which moreover also holds for the Matérn and RationalQuadratic outer kernels, cf. Eqs. (12) and (13), respectively. Figure 12 exemplifies this for the QFMNIST ($d=8$) dataset for Gaussian outer kernel function.
Figure 12: Comprehensive insight into various mechanisms of PQKs exemplified for the QFMNIST dataset with $d=8$ principal components and Gaussian outer kernel function. The central heatmap plot illustrates the $d(G|G^{\prime})$ distance measure according to Eq. (19), where each section correpsonds to a family of projected quantum circuits as defined by a given measurement operator $O^{k}$. The sections are further subdivided into the different encoding circuits in the order as they appear in Appendix B. The plot at the top details $\mathrm{Var}(F)$; cf. Eq. (20). The close-up images on the very right correspond to examples of $F$ and $F^{\prime}$ that lead to $d(G|G^{\prime})\to 0$ and $d(G|G^{\prime})\neq 0$, respectively. The heatmap visualization of $F$ can be viewed as illustrations of embeddings into the corresponding Hilbert space. The histogram of MSEs right to the central heatmap helps to identify Gram matrices that lead to good/bad performances (with bad defined as MSE $>0.1$, indicated by orange dashed line), respectively. The $w_{\mathrm{e}}$- and $\gamma$-plots on the left and on the bottom, respectively helps to understand the machanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large.
Each section in the central $d(G|G^{\prime})$ heatmap plot corresponds to a family of projected quantum circuits defined by a given measurement operator $O^{k}$ and is further subdivided into the different encoding circuits in the order as they appear in Appendix B. There are regions with finite distances between two Gram matrices and regions with vanishing differences. This can be explained considering the $\mathrm{Var}(F)$ plot at the top as well as the zoomed-in examples of the $F$ heatmaps on the very right. Vanishing $d(G|G^{\prime})$ distances correspond to Gram- and $F$-matrices that show favourable variance behavior or, in other words, lead to similar embeddings $F$ and $F^{\prime}$ in the corresponding Hilbert space (upper zoom-in plot). In contrast, finite distances (lower zoom-in plot) result from one Gram-/$F$-matrix being ill-conditioned, i.e., $\mathrm{var}(G)\to 0$ and thus $\mathrm{var}(F)\to 0$ or vice versa. Consequently, these matrices correspond to weakly performing models, while the others show better MSE scores as can be observed from the respective histogram plot on the right. Considering the $w_{\mathrm{e}}$- and $\gamma$-histograms on the left and on the bottom helps to understand the mechanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large. With this mechanism we can a posteriori explain the origin of bad performing PQK models in Figs. 4 and 8. The analogous plots for the QFMNIST($d=8$) dataset for the remaining outer kernel functions are detailed in the Appendix D.
@@ V Discussion
The study of hyperparameter importances for effective model tuning illustrates that regularization, outer kernel length scale (for PQKs) and data prescaling hyperparameters, are mostly driving the successful training of QKMs within our simulations. These results further indicate that the number of qubits and layers of the respective data encoding circuits are less important, both for FQK and PQK approaches. However, for increasing dataset complexity, in particular for the two classification tasks considered here, our findings indicate an increased importance of the number of qubits and layers. Hence, identifying and applying sufficiently complex datasets may pave the way towards a deeper understanding of quantum kernel methods. In this respect we note that it is known that quantum models can be written as partial Fourier series in the data, where the accessible frequencies are eventually determined by the number of qubits and layers of the underlying data encoding circuits [14]. Moreover, Ref. [11] outlines the Fourier representation of quantum kernels. However, it remains an interesting open question how this can be leveraged for the QKM design as well as for the identification of promising dataset properties.
Furthermore, our analysis of the model performances demonstrates that the average feature correlation to the targets can serve as a useful difficulty measure for predicting test performance scores. Interestingly, for regression tasks the particular choice of kernel method, i.e., QKRR and QSVR approaches is not really important as, on average, they are performing equally well. Similarly, our experiments across both regression and classification tasks show no significant performance differences between FQK and PQK, even though we simulate up to 15 qubits. In this range, the first effects of the limited trainability of FQKs due to the exponential concentration phenomenon [28] should appear, and one might expect evidence of better PQK performance. The absence of this observation however, may be explained by bandwidth tuning (here, we optimize $w_{\mathrm{e}}$), whereby generalizability can be established, but, as shown in Refs. [29, 37], at the expense of becoming classically tractable. To properly understand the connections between bandwidth tuning, exponential concentration and classical tractability in terms of the geometric difference metric [16] will be part of a future work. Our Spearman correlation analyses further validates the importance of classical hyperparameters, aligning with our findings from the hyperparameter importance study.
@@
Another remarkable observation from our explorations confirms Ref. [34] and demonstrates that data encoding circuits without entanglement perform well. Based on our results, we are inclined to stress this even more to that extent that encoding circuits without entanglement perform equally well or better than those with entangling gates. This holds for FQKs and PQKs as well as for QSVM and QKRR approaches in regression and classification tasks, respectively, highlighting the need to unravel the driving forces of QKM performance if “quantumness” turns out as a debatable mechanism.
@@
We note that although we have randomly initialized the trainable parameters of data encoding circuits, they resulted in comparatively good performance scores in all cases. Although the influence of tuning these variational parameters has not been investigated systematically in this study, we found in some cases that training them can improve the QKM performance, although there is no guaranty, cf. Appendix E. Hence, we recommend to carefully scrutinize this in future studies.
The survey of different data encoding strategies in terms of embedding some or all features redundantly reveals that for tasks involving datasets with lower complexity, based on FQKs, it is beneficial to redundantly encode all features on a number of qubits that is given by integer multiples of the number of features. For PQKs, redundant embedding appears to be negligible. For tasks involving datasets with higher complexity, we observed the best performances for both FQK and PQK for a number of qubits corresponding to non-integer multiples of the number of features.
Our exploration of PQK design options reveals that projected quantum circuits, outer kernel length scale parameters, regularization strengths, and data prescaling hyperparameters drive the learning capability of PQK-based QKMs synergistically. At this, we note that investigating shadow kernels or linear PQKs [16] accordingly remains open.
Beyond that, we again emphasize the need for proper and sufficiently complex datasets to obtain deeper insights into quantum kernel models. In regard to PQKs, this may promote our understanding of the impact of the choice of the measurement operator and the outer kernel. Related to this, it is worth mentioning again that although we encompass multiple different dataset families, the effective number of distinct families, the size of each dataset, and the specifics of splitting into training and test sets represents a restriction, which limits the generalizability of our results.
In conclusion we believe that this study substantially contribute to better understand what distinguishes good from bad QKMs. For the datasets we investigated in this work, we showed that in most cases it remains unresolved to clearly reveal the quantum contribution and its importance for learning. At most for PQKs we could reveal that classical and quantum parts can contribute likewise. In our view, these findings corroborate the need to tackle QML research from two sides, i.e.  identify requirements for datasets with potential quantum advantage and explore how corresponding model design should look like.
Acknowledgements.

### v3

@@ 4.3 Influence of Encoding Circuits
Figure 8: Impact of data encoding circuits on test performance scores. Results are aggregated across all datasets within the respective family and $n_{\mathrm{layers}}\in[1,8]$. The upper panel exemplifies this for the Friedman regression problem. In (a) we show results corresponding to the dataset with $d=15$ features. Here, test performance scores, as measured by MSE, are aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel (FQK/PQK). The findings in (b) represent test performance scores aggregated over all datasets in the Friedman family. The lower panel details the two curves diff dataset. In (c), we display test performances, as measured by ROC-AUC score, aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with degree $D=13$. Aggregated results across all datasets are depicted in (d).
@@
To investigate the impact of data encoding circuits on the performance of the resulting models, we aggregate results across all datasets within each family and choose $n_{\mathrm{layers}}\in[1,8]$ The results per encoding circuit are shown in Fig. 8 for the Friedman and the two curves diff datasets. The results of the other dataset families of this study are shown in Fig. A12 of the Appendix G.3. When examining the performance scores for a single dataset within a dataset family, significant performance differences among the data encoding circuits emerge. This can be seen in Figs. 8 (a) and (c). Here, the “HighDimEncodingCircuit” and the “ZZFeatureMap” perform particularly poor in regression tasks, while the “ChebyshevPQC encoding circuit” underperforms in classification tasks, both for FQK and PQK approaches (note that trends for MSE and ROC-AUC performance scores work in opposite directions). However, when considering performance scores aggregated across all datasets within a dataset family, we observe that, except for the “ChebyshevPQC” in the QSVC-FQK case and with some trend for the “ZZFeatureMap” in regression tasks, all encoding circuits perform comparably in terms of overall performance and variance (cf. Figs. 8 (b) and (d)). This supports arguments for problem-specific data encoding architecture search [2, 43].
@@
In earlier studies [8], it has been found that data encoding circuits without entanglement can perform surprisingly well. In our case, these are the “SeparableRx” and “ZFeatureMap” encoding circuits. With associated model performances obtained in Fig. 8, this raises the possibility that driving forces other than “quantumness” might be fostering the performance of QKMs. This observation underscores the need for further comprehensive and systematic investigations into the underlying factors contributing to QKM performance.
In Appendix G.4, we additionally investigate the effect of different data encoding strategies to distribute the features of a data point among the available qubits in a particular encoding circuit. Here, we find that depending on the dataset complexity it can be beneficial to redundantly encode all or merely certain features, respectively. Overall, the observed effects are, however, small compared to other factors such as adequate hyperparameter tuning.
4.4 Analysis of PQK Design Options
@@
Table 1: Semi-partial correlation analysis results for exploring the effect of the choice of the outer kernel function $\kappa$ (cf. (7)) and the measurement operator $O_{k}$ (cf. Eqs. (10), (12), (14)) in the definition of a PQK on the resulting test performance scores, while controlling for the effect of all other hyperparameters on the test performance. We investigate PQK-QSVC and PQK-QSVR models for classification and regression tasks, respectively. We compute ROC-AUC scores for evaluating classification accuracy and MSEs for assessing regression results. We employ Spearman $\rho$ rank-order correlation analysis as implemented in the pingouin [57] package. Bold values refer to statistically significant correlations, with $p$-values0.05absent0.05$\leq$ 0.05≤ 0.05, while others refer to not statistically significant correlation, which are given for the sake of completeness. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
@@
Projected quantum kernels as defined in Eq. (7) have additional degrees of freedom which can strongly influence the performance of the resulting model. Specifically, any proper outer kernel function $\kappa$ can be applied and the measurement operator $O_{k}$ can be freely selected. To the best of our knowledge, a thorough investigation of the influence of both of these design choices on the model is still lacking. [^5] Reference [16] touches upon similar aspects but has a much narrower focus on the impact of the RDM size on cross-validation accuracy and generalization ability.
[…]
@@
To quantify the influence of the respective choices on the model performance, while taking into account the corresponding impact of all other hyperparameters, we aggregate all simulation results per dataset and conduct (semi-) partial correlation analyses (cf. Appendix. B for details). The results are summarized in Tab. 1. Note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, the opposite is the case for correlations with ROC-AUC scores. An overview of the model performances used for calculating the values in the table is provided in Appendix G.5 in Figs. A19 and A20. From the table we directly realize that both $\kappa$ and the choice of $O_{k}$ for defining the respective projected quantum circuits exhibit statistically significant correlations. For the outer kernel, no universal statement can be made a priori regarding which one is generally suitable for effective QKM design—at least not within the scope of our study. The only exception is that for the classification tasks considered here, the RationalQuadratic kernel is not a good choice. Furthermore, it is worth noting that the commonly used Gaussian kernel is not automatically the best choice.
@@
Regarding a proper choice of the measurement operator $O_{k}$, Tab. 1 suggests $X_{k=1}$ to be insufficient for all regression and classification tasks considered here. In this case, we frequently observed training kernel Gram matrices entirely filled with ones, thus rendering them impractical for learning. In contrast, utilizing the $P^{1+2}$ measurement operator, cf. Eq. (14), points towards an overall improvement in model performance (within the scope of this study).
Figure 9: Comprehensive insight into various mechanisms of PQKs exemplified for the QFMNIST dataset with $d=8$ principal components and Gaussian outer kernel function. The central heatmap plot illustrates the $d(G|G^{\prime})$ distance measure according to Eq. (15), where each section correpsonds to a family of projected quantum circuits as defined by a given measurement operator $O_{k}$. The sections are further subdivided into the different encoding circuits in the order as they appear in Appendix C. The plot at the top details $\mathrm{Var}(F)$; cf. Eq. (16). The close-up images on the very right correspond to examples of $F$ and $F^{\prime}$ that lead to $d(G|G^{\prime})\to 0$ and $d(G|G^{\prime})\neq 0$, respectively. The heatmap visualization of $F$ can be viewed as illustrations of embeddings into the corresponding Hilbert space. The histogram of MSEs right to the central heatmap helps to identify Gram matrices that lead to good/bad performances (with bad defined as MSE $>0.1$, indicated by orange dashed line), respectively. The $w_{\mathrm{e}}$- and $\gamma$-plots on the left and on the bottom, respectively helps to understand the machanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large.
@@
In Ref. [8] the authors have raised the question whether the projected quantum circuit is not so much responsible for learning, but rather the subsequent outer (Gaussian) kernel applied to the features computed by Eq. (2.2). Referring to Tab. 1 per dataset, it becomes evident that both the choice of $O_{k}$ (and thus the projected quantum circuit itself), and the outer kernel $\kappa$ are crucial for achieving sufficiently good model performance. Therefore, the question of which factor is more important cannot be definitively answered based on this analysis alone. Consequently, we further list all (statistically significant) correlations of the various hyperparameters of the PQK QKMs for a given outer kernel, both among themselves and with model performance in Appendix G.5 in Tab. 2. Thereof, it again becomes clear that depending on the dataset and outer kernel function, both hyperparameters that contribute to the definition of the projected quantum circuits (i.e., $n_{\mathrm{layers}}$ and $w_{\mathrm{e}}$) and those solely associated with the outer kernel exhibit statistically significant correlations.
@@
where $|G|$ refers to the number of entries in $G$. For Gaussian outer kernels it directly follows from Eq. (8) that the projected quantum circuit contribution $F_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ becomes
$F_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=-\frac{\log[k_{\mathbf{\theta}}^{\text{PQK}}(\mathbf{x},\mathbf{x}^{\prime})]}{\gamma}\,.$ (16)
Thus, calculating $d(G|G^{\prime})$ also defines a difference measure between projected quantum circuit contributions $F$ and $F^{\prime}$, which moreover also holds for the Matérn and RationalQuadratic outer kernels, cf. Eqs. (A.1) and (19), respectively. Figure 9 exemplifies this for the QFMNIST ($d=8$) dataset for Gaussian outer kernel function. Each section in the central $d(G|G^{\prime})$ heatmap plot corresponds to a family of projected quantum circuits defined by a given measurement operator $O_{k}$ and is further subdivided into the different encoding circuits in the order as they appear in Appendix C. There are both regions with finite distances between two Gram matrices and regions with vanishing differences. This can be explained by the $\mathrm{Var}(F)$ plot at the top as well as the zoomed-in examples of the $F$ heatmaps on the very right. Vanishing $d(G|G^{\prime})$ distances correspond to Gram- and $F$-matrices that show favourable variance behavior or, in other words, lead to similar embeddings $F$ and $F^{\prime}$ in the corresponding Hilbert space (upper zoom-in plot). In contrast, finite distances (lower zoom-in plot) result from one Gram-/$F$-matrix being ill-conditioned, i.e., $\mathrm{var}(G)\to 0$ and thus $\mathrm{var}(F)\to 0$ or vice versa. Consequently, these matrices correspond to weakly performing models, while the others show better MSE scores as can be observed from the respective histogram plot on the right. Considering the $w_{\mathrm{e}}$- and $\gamma$-histograms on the left and on the bottom helps to understand the mechanisms which render these cases into the weak-performing regime. Here, $w_{\mathrm{e}}$ becomes comparatively small, while $\gamma$ becomes relatively large. With this mechanism we can a posteriori explain the origin of bad performing PQK models in Figs. 4 and 8. The analogous plots for the QFMNIST($d=8$) dataset for the remaining outer kernel functions are detailed in the Appendix G.
@@ 5 Discussion
Our findings emphasize the importance of hyperparameters such as regularization and bandwidth-tuning as well as outer kernel length scale and the choice of measurement operators for PQKs, in effective training of QKMs. Interestingly, the number of qubits and layers in data encoding circuits appear less crucial. Performance comparisons reveal no substantial differences between FQKs and PQKs. For less complex datasets, the particular choice of the kernel method is not really important as, on average, they are performing equally well. While PQKs slightly outperform classical baselines for simpler datasets, both quantum and classical models struggle as the dataset complexity increases, indicating no clear quantum advantage.
@@
Another remarkable observation significantly extends the findings of Ref. [8] and generlly reveals that circuits without entanglement perform on par with or better than those with entangling gates. We show this across various encoding circuits, using both FQKs and PQKs within QSVM and QKRR approaches and addressing both regression and classification problems. Thus, we highlight the need to unravel the driving forces of QKM performance if “quantumness” turns out as debatable mechanism.
@@
We note that although we encompass multiple different dataset families, the effective number of distinct families, the size of each dataset, and the specifics of splitting into training and test sets represents a restriction, which limits the generalizability of our results. Moreover, we randomly initialized the trainable parameters of data encoding circuits. While this resulted in comparatively good performance scores in all cases, we found in some cases that training them can improve the QKM performance, although there is no guaranty, cf. Appendix H. Hence, we recommend to carefully scrutinize this in future studies.
Our 15-qubit simulations find no clear PQK performance advantage despite expected FQK trainability limits as a consquence of exponential concentration [55], likely due to bandwidth tuning improving generalization [52, 11] but, as shown in Refs. [53, 16], at the expense of becoming classically tractable. To properly understand the connections between bandwidth tuning, exponential concentration and classical tractability in terms of the geometric difference metric [22] will be part of a future work.
In the context of our results, we believe that future research should focus on developing and applying datasets with larger complexity to gain a deeper understanding of quantum kernel methods. One idea could be to start form the Fourier representation of quantum kernels [51, 49] and investigate how this can be leveraged for the QKM design as well as for the identification of promising dataset properties. If such a dataset is found, then thoroughly investigating corresponding implications on generalization properties and exponential concentration [55] as well as classical tractability [22] of those quantum kernels would be of great interest. Another crucial aspect is to systematically explore problem-specific data encoding strategies.
6 Conclusion
This study provides valuable insights into the design of effective QKMs, offering essential guidlines for achieving good model performance. Key factors include data preprocessing, bandwidth-tuning, and careful optimization of classical hyperparameters for both FQKs and PQKs. While the number of qubits and layers of underlying data encoding circuits becomes only critical as the dataset complexity increases, the choice of outer kernel functions and measurement operators is vital for PQKs. Here, we revealed that classical and quantum parts drive learning synergistically.
Despite identifying these universal patterns, generally unraveling the true quantum contribution of QKMs remains elusive. Although we showed that more complex datasets require more expressive encodings, our findings challenge how QKMs leverage quantum-specific advantages, as encoding circuits without entanglement perform equally well or even better. With this, our work underscores the need for a dual approach to QML research: identifying datasets with potential quantum advantage and refining the corresponding model designs to fully exploit quantum capabilities.
[…]